# DGInStyle: Domain-Generalizable Semantic Segmentation with Image Diffusion Models and Stylized Semantic Control

### v2

@@ Abstract
Large, pretrained latent diffusion models (LDMs) have demonstrated an extraordinary ability to generate creative content, specialize to user data through few-shot fine-tuning, and condition their output on other modalities, such as semantic maps. However, are they usable as large-scale data generators, e.g., to improve tasks in the perception stack, like semantic segmentation? We investigate this question in the context of autonomous driving, and answer it with a resounding ”yes”. We propose an efficient data generation pipeline termed DGInStyle. First, we examine the problem of specializing a pretrained LDM to semantically-controlled generation within a narrow domain. Second, we propose a Style Swap technique to endow the rich generative prior with the learned semantic control. Third, we design a Multi-resolution Latent Fusion technique to overcome the bias of LDMs towards dominant objects. Using DGInStyle, we generate a diverse dataset of street scenes, train a domain-agnostic semantic segmentation model on it, and evaluate the model on multiple popular autonomous driving datasets. Our approach consistently increases the performance of several domain generalization methods compared to the previous state-of-the-art methods. Source code and dataset are available at dginstyle.github.io.
@@ 1 Introduction
Semantic segmentation requires large annotated datasets for supervised training. While manual annotation is costly and time-consuming [11, 55], synthetic datasets offer a cost-effective solution. However, these datasets face a domain gap [16], leading to poor performance when networks trained on the source domain are applied to the target domain. When the characteristics of the target domain are known, the domain gap can be addressed with Domain Adaptation techniques [16, 26]. A more challenging, arguably equally important setting is Domain Generalization (DG) [76, 28, 14], where a model is deployed in a new environment without knowing the target domain except for its general context (such as “autonomous driving”).
In the DG semantic segmentation literature, the role of the prior domain is often overlooked. In end-to-end pipelines, that prior typically remains implicit; for instance, it could stem from pretrained backbone weights used in most segmentation DG methods (often from ImageNet [12]) or loss functions that depend on feature space distances [28, 76]. Therefore, we take a closer look at the prior domain and study how we can utilize the rich prior that emerges in modern foundational models trained on internet-scale datasets [57] to improve domain generalization of semantic segmentation. To this end, we design DGInStyle, a novel data generation pipeline with a pretrained foundational text-to-image LDM [50] at its core, fine-tuned with data from the source domain, with conditioning on the associated dense label maps. Such a pipeline can automatically generate images with characteristics of the prior domain and equipped with pixel-aligned label maps (Fig. 1). Armed with such a pipeline, we approach DG differently from other methods by focusing on synthesizing data instead of model architectures or training techniques. The idea is that a model trained on such data will offer improved domain generalization, drawing on the prior knowledge embedded in the LDM.
@@
This comes with two new challenges: The LDM needs to learn to produce images that match semantic masks from the labeled source domain while avoiding overfitting to its style. Additionally, the generated images need precise alignment with segmentation masks, even for very small instances. Our Contributions address these issues: First we propose a novel Style Swap technique inspired by modern fine-tuning and semantic style control mechanisms, to achieve the necessary level of control and diversity over the outputs. It is based on the novel finding that the semantic control and the underlying (stylized) diffusion model can be decoupled and swapped. This enables our simple yet efficient Style Swap, which allows to learn dense semantic control on the source domain while removing the undesired source domain style. Second, we present a novel Multi-Resolution Latent Fusion technique, which helps us to go beyond the limited resolution of the LDM generator. It is an essential step to achieve conditioned generation of small instances, which is crucial for learning semantic segmentation on generated data. Without it, a segmentation model trained on it would struggle with inconsistent generated images and segmentation masks. Lastly, we use the resulting generative pipeline to create a diversified dataset to train semantic segmentation networks, including methods to mitigate the impact of domain shifts. Due to its complementary design, DGInStyle achieves major performance improvements when combined with existing DG methods. In particular, it significantly boosts the state-of-the-art domain generalization in autonomous driving, as shown in Fig. 2.
@@ 2 Related Work
2.1 Generative Models
Early advancements in deep learning techniques led to a surge in deep generative models, namely GANs and VAEs [19, 34, 47, 6]. While GANs exhibited training challenges such as instability and mode collapse [5], VAEs struggled with output quality and artifacts.
@@ 2.1 Generative Models
2.2 Dataset Generation
The pioneering work DatasetGAN [74] automatically generates labeled images by manipulating feature maps from StyleGAN and outperforms semi-supervised baselines in object-part segmentation. Recent techniques have utilized state-of-the-art DMs to create training data for downstream tasks such as image classification [22, 60, 15, 56, 1, 36], object detection [73, 8, 64], semantic segmentation [62, 35, 45, 68, 18, 63].
@@ 2.2 Dataset Generation
2.3 Domain Generalization
Unsupervised Domain Adaptation (UDA) [16] focuses on learning to perform a task in a target domain through supervised learning on labeled data from a similar source domain. Only unannotated data from the target domain is available in this setting. This task received much attention due to the simplicity of the formulation; several approaches [27, 53] were proposed to efficiently bridge the domain gap.
@@ 2.3 Domain Generalization
2.4 Diffusion Models for Domain Generalization
Beyond the aforementioned approaches, recent works have explored the use of diffusion models for domain generalization in semantic segmentation. Gong et al. [18] investigate how well diffusion-pretrained representations generalize to new domains and introduce a prompt randomization strategy to enhance cross-domain performance. DatasetDM [62] presents a generic dataset generation model capable of producing diverse images and corresponding annotations using diffusion models. These methods implement grounded generation by training a segmentation decoder to achieve image-mask alignment. Our approach takes a different semantic guidance route, exhibiting higher controllability and generating consistent image-label pairs that qualify as training datasets.
@@ 3 Methods
Figure 3: Overview of our proposed Style Swap technique. ControlNet learns segmentation-conditioned image generation on the source domain. To avoid that ControlNet also steers the generated style, it is trained on top of a source domain fine-tuned LDM. So, the source domain LDM can be replaced with the original LDM to restore its rich style prior. As discussed in Sec. 4, this technique leads to state-of-the-art results in domain generalization for semantic segmentation.
Figure 4: ControlNet learns the source domain style. This effect hinders varied data generation for domain generalization. The proposed Style Swap preserves the diversity of the prior domain.
@@ 3.1 Label Conditioned Image Generation
Figure 5: Illustration of MRLF module. We generate a first-pass image $I$ using low-resolution conditioning. In the subsequent high-resolution pass, we partition the canvas into overlapping tiles at each generation step, concurrently apply denoising to each with its respective conditioning, and fuse them with a tile diffusion technique. Finally, we preserve the quality of large objects in the mask $\mathrm{M}$ with inpainting conditioned on the first pass image. The color gradient  Refer to caption   represents the path from noise to clean data.
(a) Car, Rider…
(b) Foggy
(c) Rainy
(d) Snowy
@@ 3.2 Preserving Style Prior with Style Swap
As a first step of our Style Swap technique, we fine-tune the base LDM’s U-Net $\mathbf{U}^{\mathbf{P}}$ encapsulating the prior domain with the Dreambooth protocol [52] using all images in the source domain. The resulting U-Net is denoted as $\mathbf{U}^{\mathbf{S}}$. Second, we use $\mathbf{U}^{\mathbf{S}}$ as the base model instead of $\mathbf{U}^{\mathbf{P}}$ to initialize ControlNet. The idea is to allow $\mathbf{U}^{\mathbf{S}}$ to absorb the domain style and let the ControlNet focus primarily on the task-specific yet style-agnostic layout control, thereby reducing the domain style bleeding into its weights. Finally, in the third step, we perform inference with the trained ControlNet, except that we switch the base LDM generator to $\mathbf{U}^{\text{P}}$ while keeping the ControlNet trained for $\mathbf{U^{\text{S}}}$. This allows us to apply semantic control learned from the source domain to the original LDM. The overall procedure endows the original LDM with task-specific semantic control, allowing us to generate diverse images adhering to the semantic segmentation masks (Fig. 4 (d)).
@@ 3.3 Style Prompting
Text prompting is a powerful technique for style mining. To better guide ControlNet generation, we concatenate unique class names present in the semantic mask into a list and pass it to the text encoder. We further enrich the diversity of the generated data by fusing randomized task-specific qualifiers into the text conditioning. For the autonomous driving segmentation, we use a range of adverse weather conditions (e.g., foggy, snowy, rainy, overcast, and night scenarios). An example of a text prompt can be: A city street scene photo with car, road, sky, rider, bicycle, vegetation, building in foggy weather. This approach, especially when integrated with the Style Swap technique, allows us to produce images with precise semantic layouts and varied styles, as shown in Fig. 6.
@@ 3.4 Multi-Resolution Latent Fusion
While ControlNet effectively integrates condition masks into the generation process, it struggles with generating small objects due to the low-resolution latent space. We propose a two-stage Multi-Resolution Latent Fusion pipeline to improve the adherence to semantic masks in the generated dataset. During the first low-resolution pass (Fig. 5, bottom-left), we perform a regular ControlNet generation at the original LDM resolution. This generation serves as a reference for the second, high-resolution generation pass. Therein, we keep the large segments generated initially and refine everything else. To overcome the problem of low latent space resolution problem, we perform the second pass in the upsampled latent space, followed by downsizing the generated image to the original size.
@@
Such a controlled generation in the upsampled space overcomes the low-resolution bias of the pretrained LDM and results in higher-quality small objects. Nevertheless, this procedure alone leads to a noticeable degradation of large objects due to the reduced field of view of the denoiser. We address this shortcoming by taking large objects from the first low-resolution pass and fusing them into the high-resolution pass using the Latent Inpainting Diffusion technique discussed below.
@@
As a result, our multi-resolution latent fusion scheme overcomes the resolution-specific limitations of the LDM. It unlocks controlled arbitrary resolution generation through processing tiles and, at the same time, preserves trusted regions with the latent inpainting diffusion scheme.
@@ 4.2 Implementation Details
We first conduct DreamBooth [52] fine-tuning using GTA images to obtain $\mathbf{U}^{\text{S}}$. The images are randomly resized and cropped to a resolution of $512{\times}512$. The fine-tuning takes 10k iterations with a constant learning rate of $2{\times}10^{-6}$.
The ControlNet [72] training is initialized with the source style $\mathbf{U}^{\text{S}}$. For input conditions, we use one-hot encoded GTA segmentation masks and crop them to the size of $512{\times}512$. These crops are guided by input text containing semantic classes in each crop.
During ControlNet inference, we perform the Style Swap as discussed in  LABEL:{sec:method_stylization} and integrate the multi-resolution latent fusion module with $s{=}2$. Our tiling strategy uses a 16-pixel stride between neighbor crops. We use $T{=}0.01$ for rare class sampling probability. The constructed dataset comprises an equal mix of images with basic text inputs and those with randomized adverse weather prompts. Extra examples are shown in the supplement.
To assess the efficacy of our DGInStyle, we train a semantic segmentation model on a combination of the GTA source dataset and our generated dataset. Specifically, we generate $N_{g}=6000$ images and select $N_{s}=6000$ images based on the rare class criteria. The training is performed under the aligned domain generalization framework as detailed in [28].
4.3 Comparison with State-of-the-Art DG
@@
Figure 7: Comparison of the class-wise IoU averaged over the five datasets using DAFormer with and without our dataset integration. The color visualizes the difference to the first row.
@@
Qualitative Analysis. Fig. 8 provides visual examples of semantic segmentation results obtained from HRDA trained with or without the use of DGInStyle. It shows that our generated dataset improves the predictions, even for difficult classes (e.g., truck and pole).
@@ 4.4 Ablation Studies
We conduct ablation studies to evaluate the contribution of each component of our method. The results are shown in Tab. 3. All models are based on the DAFormer [26] framework but trained on datasets generated under varying conditions. We observe that incorporating Multi-Resolution Latent Fusion (MRLF) enhances the generation of small objects in our dataset, boosting the segmentation model’s performance by +0.96 on average of the five datasets. As a vital part of the style diversification module, the Style Swap technique significantly improves model performance by another +1.57, demonstrating the effectiveness of utilizing the prior domain to generate diverse samples. The Style Prompts module further elevates the model performance by +0.32, especially in adverse weather scenarios [55, 54]. Combined with Rare Class Generation (RCG), which adds another +0.51, our complete data generation pipeline achieves an average mIoU of 46.67% over the five real-world datasets.
We additionally present the ablations by excluding each component during the dataset generation to evaluate their role in the combined framework. As seen in Tab. 3, removing the Style Swap component significantly degrades performance, underscoring its effectiveness in leveraging prior knowledge to diversify the generated data. Similarly, removing other components also leads to a decline in the model’s performance, which reveals that each component adds value to our data generation pipeline.
Figure 9: Qualitative examples of MRLF. (a) When zooming in on the mask crop, which contains small objects such as cars and traffic poles, the initial generation fails to generate recognizable contents. (b) This is addressed by conducting Controlled Tiled MultiDiffusion, which enhances the generation quality in fine-grained granularity. However, it can lead to artifacts of large objects. (c) When adding Latent Inpainting Diffusion, the generated image not only improves the local details but also reduces artifacts in large objects.
@@
53.07 45.19
54.05 45.60
54.25 46.67
Table 4: MRLF Ablation. Ablation studies on multi-resolution components with Controlled Tiled MultiDiffusion (CTMD) and Latent Inpainting Diffusion (LID). Numbers are reported in mIoU.

### v3

@@ Abstract
Large, pretrained latent diffusion models (LDMs) have demonstrated an extraordinary ability to generate creative content, specialize to user data through few-shot fine-tuning, and condition their output on other modalities, such as semantic maps. However, are they usable as large-scale data generators, e.g., to improve tasks in the perception stack, like semantic segmentation? We investigate this question in the context of autonomous driving, and answer it with a resounding "yes". We propose an efficient data generation pipeline termed DGInStyle. First, we examine the problem of specializing a pretrained LDM to semantically-controlled generation within a narrow domain. Second, we propose a Style Swap technique to endow the rich generative prior with the learned semantic control. Third, we design a Multi-resolution Latent Fusion technique to overcome the bias of LDMs towards dominant objects. Using DGInStyle, we generate a diverse dataset of street scenes, train a domain-agnostic semantic segmentation model on it, and evaluate the model on multiple popular autonomous driving datasets. Our approach consistently increases the performance of several domain generalization methods compared to the previous state-of-the-art methods. The source code and the generated dataset are available at dginstyle.github.io.
Keywords:
Semantic Domain Generalization Image Latent Diffusion
@@ 1 Introduction
Segmenting images into semantically defined categories requires large annotated datasets of images and associated label maps, as a basis for supervised training. Manual annotation for obtaining those label maps is time-consuming and expensive [11, 56], which is where image generation comes into play. Synthetic datasets are annotated by construction and therefore cheap to collect, but they invariably suffer from a domain gap [16], meaning that a network trained on such data (the source domain) will perform poorly on the real images of interest (the target domain). When the characteristics of the target domain are known in advance through (labeled or unlabeled) samples, the domain gap can be addressed with Domain Adaptation techniques [16, 26]. A more challenging, arguably equally important setting is Domain Generalization (DG) [77, 28, 14], where a model is deployed in a new environment without the chance to first collect data and adapt. I.e., the target domain is unknown except for the high-level application context (such as “autonomous driving”).
In the DG semantic segmentation literature, the role of the prior domain is often overlooked. In end-to-end pipelines, that prior typically remains implicit; for instance, it could stem from pretrained backbone weights used in most segmentation DG methods (often from ImageNet [12]) or loss functions that depend on feature space distances [28, 77]. Therefore, we take a closer look at the prior domain and study how we can utilize the rich prior that emerges in modern foundational models trained on internet-scale datasets [58] to improve domain generalization of semantic segmentation.
To this end, we design DGInStyle, a novel data generation pipeline with a pretrained foundational text-to-image LDM [51] at its core, fine-tuned with data from the source domain, with conditioning on the associated dense label maps. Such a pipeline can automatically generate images with characteristics of the prior domain and equipped with pixel-aligned label maps (Fig. 1). Armed with such a pipeline, we approach DG differently from other methods by focusing on synthesizing data instead of model architectures or training techniques. The idea is that a model trained on such data will offer improved domain generalization, drawing on the prior knowledge embedded in the LDM.
This comes with two important new challenges: The LDM needs to learn how to produce images that match semantic segmentation masks. This can only be learned on the labeled source domain. However, during the process, the LDM must not overfit to the source domain style. Additionally, the generated images must exactly align with segmentation masks, even for very small instances.
@@
Therefore, several fundamental modifications are necessary to turn an off-the-shelf LDM [51] into a data generation pipeline for domain-generalizable semantic segmentation, which would otherwise suffer from source domain style bleeding and ignoring small instances. Our Contributions address these issues: First we propose a novel Style Swap technique inspired by modern fine-tuning and semantic style control mechanisms, to achieve the necessary level of control and diversity over the outputs. It is based on the novel finding that the semantic control and the underlying (stylized) diffusion model can be decoupled and swapped. This enables our simple yet efficient Style Swap, which allows to learn dense semantic control on the source domain while removing the undesired source domain style. Second, we present a novel Multi-Resolution Latent Fusion technique, which helps us to go beyond the limited resolution of the LDM generator. It is an essential step to achieve conditioned generation of small instances, which is crucial for learning semantic segmentation on generated data. Without it, a segmentation model trained on it would struggle with inconsistent generated images and segmentation masks. Lastly, we use the resulting generative pipeline to create a diversified dataset to train semantic segmentation networks, including methods to mitigate the impact of domain shifts. Due to its complementary design, DGInStyle achieves major performance improvements when combined with existing DG methods. In particular, it significantly boosts the state-of-the-art domain generalization in autonomous driving, as shown in Fig. 2.
@@ 2 Related Work
Generative Models. Early advancements in deep learning techniques led to a surge in deep generative models, namely GANs and VAEs [19, 34, 48, 7]. While GANs exhibited training challenges such as instability and mode collapse [5], VAEs struggled with output quality and artifacts.
@@
Dataset Generation. The pioneering work DatasetGAN [75] automatically generates labeled images by manipulating feature maps from StyleGAN and outperforms semi-supervised baselines in object-part segmentation. Recent techniques have utilized state-of-the-art DMs to create training data for downstream tasks such as image classification [22, 61, 15, 57, 1, 36], object detection [74, 8, 65], semantic segmentation [63, 35, 44, 69, 18, 64].
@@
Domain Generalization. Unsupervised Domain Adaptation (UDA) [16] focuses on learning to perform a task in a target domain through supervised learning on labeled data from a similar source domain. Only unannotated data from the target domain is available in this setting. This task received much attention due to the simplicity of the formulation; several approaches [27, 54] were proposed to efficiently bridge the domain gap.
@@
Diffusion Models for Domain Generalization. Beyond the aforementioned approaches, recent works have explored the use of diffusion models for domain generalization in semantic segmentation. Gong et al. [18] investigate how well diffusion-pretrained representations generalize to new domains and introduce a prompt randomization strategy to enhance cross-domain performance. DatasetDM [63] presents a generic dataset generation model capable of producing diverse images and corresponding annotations using diffusion models. These methods implement grounded generation by training a segmentation decoder to achieve image-mask alignment. Our approach takes a different semantic guidance route, exhibiting higher controllability and generating consistent image-label pairs that qualify as training datasets.
@@ 3.2 Preserving Style Prior with Style Swap
(a) Source Mask (b) Source Image (c) Gen. w/o Swap (d) Gen. w/ Swap
Figure 3: ControlNet learns the source domain style. This effect hinders varied data generation for domain generalization. Our Style Swap mitigates the effect and preserves the style prior.
(a) Car, Rider… (b) Foggy (c) Rainy (d) Snowy
Figure 5: Overview of our proposed Style Swap technique. ControlNet learns segmentation-conditioned image generation on the source domain. To avoid ControlNet steering the generated style, it is trained on top of a source domain fine-tuned LDM. Later, this source domain LDM can be replaced with the original LDM to restore the rich style prior. As discussed in Sec. 4, this technique leads to state-of-the-art results in domain generalization for semantic segmentation.
@@
As a first step of our Style Swap technique, we fine-tune the base LDM’s U-Net $\mathbf{U}^{\mathcal{P}}$ encapsulating the prior domain $\mathcal{P}$ with the Dreambooth protocol [53] using all images in the source domain $\mathcal{S}$. The resulting U-Net is denoted as $\mathbf{U}^{\mathcal{S}}$. Second, we use $\mathbf{U}^{\mathcal{S}}$ as the base model instead of $\mathbf{U}^{\mathcal{P}}$ to initialize ControlNet. The idea is to allow $\mathbf{U}^{\mathcal{S}}$ to absorb the domain style and let the ControlNet focus primarily on the task-specific yet style-agnostic layout control, thereby reducing the domain style bleeding into its weights. Finally, in the third step, we perform inference with the trained ControlNet, except that we switch the base LDM generator to $\mathbf{U}^{\mathcal{P}}$ while keeping the ControlNet trained for $\mathbf{U^{\mathcal{S}}}$. This enables us to apply semantic control learned from the source domain to the original LDM. The overall procedure endows the original LDM with task-specific semantic control, allowing us to generate diverse images adhering to the semantic segmentation masks. This result is shown in Fig. 4 (d).
@@ 3.3 Style Prompting
Text prompting is a powerful technique for style mining. To better guide ControlNet generation, we concatenate unique class names present in the semantic mask into a list and pass it to the text encoder. We further enrich the diversity of the generated data by fusing randomized task-specific qualifiers into the text conditioning. These can be obtained from the task definition with a query to a domain expert or an LLM, and sometimes are known in advance, e.g., from the source data simulator, such as GTA [49]. For the autonomous driving segmentation, we use a range of adverse weather conditions (e.g., foggy, snowy, rainy, overcast, and night scenarios). An example text prompt can be: A city street scene photo with car, road, sky, rider, bicycle, vegetation, building, in foggy weather. This approach, especially when integrated with the Style Swap technique, allows producing images with semantic control and varied natural styles from the prior domain $\mathcal{P}$, as shown in Fig. 4.
@@ 3.4 Multi-Resolution Latent Fusion
While ControlNet effectively integrates condition masks into the generation process, it struggles with generating small objects due to the low-resolution latent space. We propose a two-stage Multi-Resolution Latent Fusion pipeline to improve the adherence to semantic masks in the generated dataset. During the first low-resolution pass (Fig. 6, bottom-left), we perform a regular ControlNet generation at the original LDM resolution. This generation serves as a reference for the second, high-resolution generation pass. Therein, we keep the large segments generated initially and refine everything else. To overcome the problem of low resolution of the latent space, we perform the second pass in the upsampled latent space, followed by downsizing the generated image to the original size.
Figure 6: MRLF module. We generate a first-pass image $I$ using low-resolution conditioning. In the subsequent high-resolution pass, we partition the canvas into overlapping tiles at each generation step, concurrently apply denoising to each with its respective conditioning, and fuse them with a tile diffusion technique. Finally, we preserve the quality of large objects in the mask $\mathrm{M}$ with inpainting conditioned on the first pass image. The color gradient  Refer to caption   represents the path from noise to clean data.
@@
Such a controlled generation in the upsampled space overcomes the low-resolution bias of the pretrained LDM and results in higher-quality small objects. Nevertheless, this procedure alone leads to a noticeable degradation of large objects due to the reduced field of view of the denoiser. We address this shortcoming by taking large objects from the first low-resolution pass and fusing them into the high-resolution pass using the Latent Inpainting Diffusion technique.
@@
As a result, our multi-resolution latent fusion scheme overcomes the resolution-specific limitations of the LDM. It unlocks controlled arbitrary-resolution generation through processing tiles. At the same time, it preserves trusted regions with the latent inpainting diffusion scheme.
@@ 4.2 Implementation Details
Our model is based on Stable Diffusion 1.5 [51] and requires a single consumer-grade GPU for training. We first conduct DreamBooth [53] fine-tuning using GTA images to obtain $\mathbf{U}^{\mathcal{S}}$. The images are randomly resized and cropped to a resolution of $512{\times}512$. The fine-tuning takes 10k iterations with a constant learning rate of $2{\times}10^{-6}$.
The ControlNet [73] training is initialized with the source style $\mathbf{U}^{\mathcal{S}}$. For input conditions, we use one-hot encoded GTA segmentation masks and crop them to the size of $512{\times}512$. These crops are guided by input text containing semantic classes in each crop. During ControlNet inference, we perform the Style Swap as discussed in  LABEL:{sec:method_stylization} and integrate the multi-resolution latent fusion module with $s{=}2$. Our tiling strategy uses a 16-pixel stride between neighbor crops. We use $T{=}0.01$ for rare class sampling probability. The constructed dataset comprises an equal mix of images with basic text inputs and those with randomized adverse weather prompts. Extra examples are shown in the supplement.
To assess the efficacy of our DGInStyle, we train a semantic segmentation model on a combination of the GTA source dataset and our generated dataset. Specifically, we generate $N_{\mathcal{G}}=6000$ images and select $N_{\mathcal{S}}=6000$ images based on the rare class criteria. The training is performed under the aligned domain generalization framework as detailed in [28].
4.3 Comparison with State-of-the-Art DG
@@
Figure 7: Class-wise IoU averaged over the five datasets using DAFormer with and without our dataset integration. The color visualizes the difference to the first row.
@@
Qualitative Analysis. Fig. 8 provides visual examples of semantic segmentation results obtained from HRDA trained with or without the use of DGInStyle. It shows that our generated dataset improves the predictions, even for difficult classes (e.g., truck and pole) and lighting conditions (e.g., day and night).
@@ 4.4 Ablation Studies
Semantic Mask (a) w/o MRLF (b) w/ CTMD (c) w/ LID + CTMD
Figure 9: Qualitative examples of MRLF. (a) When zooming in on the mask crop, which contains small objects such as cars and traffic poles, the initial generation fails to create recognizable content for these instances. (b) This is addressed by conducting Controlled Tiled MultiDiffusion, which enhances the generation quality of fine details. However, it can lead to artifacts of large objects. (c) When adding Latent Inpainting Diffusion, the generated image not only improves the local details but also reduces artifacts in large objects.
We conduct ablation studies to evaluate the contribution of each component of our method. The results are shown in Tab. 3. All models are based on the DAFormer [26] framework but trained on datasets generated under varying conditions. We observe that incorporating Multi-Resolution Latent Fusion (MRLF) enhances the generation of small objects in our dataset, boosting the segmentation model’s performance by +0.96 on average of the five datasets. As a vital part of the style diversification module, the Style Swap technique significantly improves model performance by another +1.57, demonstrating the effectiveness of utilizing the prior domain to generate diverse samples. The Style Prompts module further elevates the model performance by +0.32, especially in adverse weather scenarios [56, 55]. Combined with Rare Class Generation (RCG), which adds another +0.31, our complete data generation pipeline achieves an average mIoU of 46.47% over the five real-world datasets.
We additionally present the ablations by excluding each component during the dataset generation to evaluate their role in the combined framework. Tab. 3 shows that removing the Style Swap component significantly degrades performance, underscoring its effectiveness in leveraging prior knowledge to diversify the generated data. Similarly, removing other components also leads to a decline in the model’s performance, which reveals that each component adds value to our data generation pipeline.
Table 4: MRLF Ablation. Ablation studies on multi-resolution components with Controlled Tiled MultiDiffusion (CTMD) and Latent Inpainting Diffusion (LID). Numbers are reported in mIoU (higher is better).
@@
     ✘      ✘      53.07      45.19
     ✔      ✘      54.05      45.60
     ✔      ✔      54.25      46.47